Initially a flask contains nitrogen and hydrogen with a partial pressure of 0.3 atmospheres and 0.8 atmospheres. What will be the total pressure and partial pressure of nitrogen, hydrogen and ammonia after the following reaction is complete.

N2 +3H2--> 2NH3
Answer #1
[N2] [H2] [NH3]
Initial 0.3 0.8 0
Change -x -3x 2x
Final 0.3-x 0.8-3x 2x

Hence [H2] will be the limiting reagent, so the value of x will be 0.8/3

Hence, pressure of NH3 will be 2 * 0.8/3 = 1.6/3 = 0.5333 atm

Pressure of [H2] = 0

Pressure of [N2] = 0.3 - 0.8/3 = 0.1/3 = 0.0333 atm

Total Pressure = [NH3] + [N2] = 1.6/3 + 0.1/3 = 1.7/3 = 0.5667

Partial Pressure of N2 = 0.0333 atm

Partial Pressure of NH3 = 0.5333 atm